"Cost function expresses the relationship between the cost and its determinants." Discuss this statement giving examples from any firm of your choice.
1
Expert's answer
2014-05-26T12:36:38-0400
Relation between the volume of output and the minimum necessary for its production cost is called cost function.
Let's the price of labor, i.e. the amount of money, necessary to pay for theuse of the employee for a certain time , be defined as rL, and the cost of capital - the amount of money paid for the use of the means of production for some time, - rK. Then the total cost to produce a certain number of products will be: TC = rL*L + rK*K.

Given prices of production factors the cost is determined by the value of the minimum required for output volumes of labor and capital, i.e. technology represented by a production function Q = Q (L, K). Therefore, L = L
(Q), K = K (Q), and consequently, TC = TC (Q).

The cost of production depends on many factors and these factors vary from one firm to another firm in the same industry or
from one industry to another industry. The main determinants of a cost function are:

a) plant size
b) output level
c) prices of inputs used in production,
d) nature of technology
e) managerial efficiency

It costs NIKE $5 to make a pair of sneakers...
NIKE shoe production in the Third World: The facts:

-Indonesian workers make $2.46 a day
-10,000 Indonesians went on strike to protest wages that are below subsistence level.
-"If I don't work overtime, I can't survive," says Baltazar at PT Hasi Nike factory in Jakarta. He works an average of 40 overtime hours a week.
-Vietnamese workers make $l.60 a day
-1,300 workers at the Sam Yang factory went on strike to demand a one cent per hour raise in wages. Other issues include excessive and illegal overtime and compensation for working with hazardous material.
-Chinese workers make $1.75 a day
-There is no minimum wage in China and when abuses are discovered, the wholefactory disappears. "The supervisors will get nervous and move the work to another province. It's impossible to monitor factory conditions," says Asia Monitor Resource Center in Hong Kong.